OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27 Question 10 (Multiple Choice Worth 1 points) Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (anonymous):

-9x + 25 -40x + 31

OpenStudy (perl):

Lets look at the first one Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

OpenStudy (anonymous):

okayy

OpenStudy (perl):

f ( g(x)) = f( -x + 3) = 9(-x + 3) - 2 = -9x + 9*3 -2 = -9x + 27 -2 =-9x + 25

OpenStudy (anonymous):

thanks @perl and @Navami

OpenStudy (perl):

Your welcome :)